Find the equations of the tangents to the graph of y=x+1/x that are parallel to y+2x=0




1
Expert's answer
2022-03-18T15:15:23-0400

Slope of the line y+2x=0: m=-2.

Slope of the tangent: "y'=1-\frac{1}{x^2}."

We have: "y'=m" or "1-\frac{1}{x^2}=-2."

So, "x=\frac{1}{\sqrt{3}},\;x=-\frac{1}{\sqrt{3}}."

"y(\frac{1}{\sqrt{3}})=\frac{1}{\sqrt{3}}+\sqrt{3}."

"y(-\frac{1}{\sqrt{3}})=-\frac{1}{\sqrt{3}}-\sqrt{3}."

Equations of the tangent lines:

"y-\frac{1}{\sqrt{3}}-\sqrt{3}=-2(x-\frac{1}{\sqrt{3}})" or "y=-2x+2\sqrt{3}" .

"y+\frac{1}{\sqrt{3}}+\sqrt{3}=-2(x+\frac{1}{\sqrt{3}})" or "y=-2x-2\sqrt{3}."